a) Calculate the accumulated value after 6 month of an investment of kshs 1,000 at time t=0 at the following rate of interest

i) A force of interest of 0.05%

ii) A rate of interest of 5% per annum

convertible monthly

iii) An effective rate of interest of 5% per

annum

b) Given an investment of kshs1,000,000. Find the accumulation after 5years using:

I) simple discount of 8% per annum

II) compound discount of 8% per annum

III) compound interest of 8% per annum


1
Expert's answer
2021-01-25T03:53:36-0500

a)

i) find the formula:

"PV\times e^{y\timesδ} = FV"

PV=1000

δ=0.0005

y=0.5

"1000\times e^{0.5\times0.0005} = 1000.25"

ii)find the formula:

"FV = PV(1 + i)^\u207f"

PV=1000

i=0.00417

"i=\frac{iy}{12}=\frac{0.05}{12}=0.00417"

"FV = 1000(1 + 0.00417)^6=1025.28"

iii)let's convert the effective bid to the nominal one

find using the formula:

"re=(1+\frac{i}{n})^n-1"

"0.05=(1+\frac{i}{12})^{12}-1"

i=0.0512

"FV = PV(1 + i)^\u207f=1000(1+0.0512)^6=1349.31"

b)

i) find the formula:

"FV=\frac{PV}{(1-r\times n)}"

PV=1000 000

r=0.08

n=5

"FV=\frac{1 000 000}{(1-0.8\times 5)}=1 041 666.67"

ii)find the formula:

"FV=\frac{PV}{(1-r)^n}"

PV= 1000 000

r=0.08

n=5

"FV=\frac{1 000 000}{(1-0.08)^5}=1 517 262.99"

iii)find the formula:

"FV = PV(1 + i)^\u207f"

PV= 1000 000

i=0.08

n=5

"FV = 1 000 000(1 + 0.08)^5=1 469 328.08"